OpenStudy (anonymous):

Zach ran t times around the track at school, which is one-fourth mile. He ran a total of seven-twelfths mile. Solve for t. seven forty-eighths two and one-third three-sevenths one-third

OpenStudy (anonymous):

Well 1/4 of a mile and 7/12's. 1/4 Of a mile is 3/12. So then how many times does 3/12 go into 7/12?

OpenStudy (anonymous):

My wording was a bit off sorry :p

OpenStudy (anonymous):

I think it is 2 and 1/3

OpenStudy (anonymous):

its oki thank you very much for helping me :)

OpenStudy (triciaal):

1/4* t = 7/12 to solve for t multiply each side by 4 t = 28/12 = 2 and 4/12 = as above 2 and 1/3

OpenStudy (anonymous):

thank you for explaning the answer :)

OpenStudy (triciaal):

you are welcome.